the volume of a sphere is given by v

The volume of a sphere is given by V = (4/3)pi r^3
Answer #1

Volume = 4/3 *pi*r^3

r = 925 pm = 925*10^-12 M

So, V = (4/3)*(pi*(925*10^-12)^3

= 3.32*10^-27 m3